the total price of two items is 140 $. If the price of one item is increased by 20percent and the price of the other item is decreased by 25 percent the average price of the two items is 70.50. What is the original price of the more expensive item.
Answer by mananth(16946)

the total price of two items is 140 $. If the price of one item is increased by 20percent and the price of the other item is decreased by 25 percent the average price of the two items is 70.50. What is the original price of the more expensive item.
One item cost be x
and other be y
x+y = 140..............(1)

20% = 0.2
increased by 0.2
1+0.2=1.2
(1.2x 00.75y)/2 = 70.50
1.2x+0.75y=141
multiply by 100
120x+75y=14100.............(2)
sole (1) & (2)
x= cost of item I
y= cost of item II

1.00 x + 1.00 y = 140.00 .............1
Total value
120.00 x + 75.00 y = 14,100.00 .............2
Eliminate y
multiply (1)by -75.00
Multiply (2) by 1.00
-75.00 x -75.00 y = -10,500.00
120.00 x + 75.00 y = 14,100.00
Add the two equations
45.00 x = 3,600.00
/ 45.00
x = 80.00
plug value of x in (1)
1.00 x + 1.00 y = 140.00
80.00 + y = 140.00
y = 140.00 -80
y = 60.00
y = 60.00
x= 80.00 cost of item I
y= 60.00 cost of item II
m.ananth@hotmail.ca